OpenStudy (anonymous):

\[cos3x - cosx + sin4x =0\] Find general solutions.

OpenStudy (anonymous):

sin4x = sin2(2x)

OpenStudy (anonymous):

=2 sin2x cos2x

OpenStudy (anonymous):

let me try by this...

OpenStudy (anonymous):

looks like your way would work; but it might get ugly..

OpenStudy (anonymous):

i prefer not using wolfram

OpenStudy (anonymous):

use \[\cos p-\cos q\] identity for \[\cos 3x-\cos x\]

OpenStudy (anonymous):

a substitution?

OpenStudy (anonymous):

nope \[\cos p-\cos q=-2 \sin \frac{p+q}{2}\sin \frac{p-q}{2}\]

OpenStudy (anonymous):

is that a trig identity?

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

well havent seen that identity before..i dont think it is needed..

OpenStudy (anonymous):

thats a very important identity...try to keep it in your mind \[\cos(A+B)=\cos A \cos B-\sin A \sin B\]\[\cos(A-B)=\cos A \cos B+\sin A \sin B\]subtract them \[\cos(A+B)-\cos(A-B)=-2\sin A \sin B\] let \(A+B=p\) and \(A-B=q\) to get that identity

OpenStudy (anonymous):

ohh..i never knew that you can manipulate the compound-angle formula.. thanks; i will try it.

OpenStudy (anonymous):

:)